OpenStudy (anonymous):

sin theta=-8/17 Find cos2theta

Parth (parthkohli):

Do you mean \(\cos^2\theta\)?

OpenStudy (mimi_x3):

Double angle formulae..

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sin^2\theta + \cos^2\theta = 1 }\)

Parth (parthkohli):

If it is supposed to be \(\sin^2\theta\) ^

OpenStudy (anonymous):

when pi is less than theta which is less than three pi divided by two

OpenStudy (mimi_x3):

\(cos2x=cos^{2}x-sin^{2}x => 2cos^{2}x-1 => 1-sin^{2}x\)

OpenStudy (anonymous):

No, ParthKohli, its just multiplied by two. I dont understand :(

OpenStudy (mimi_x3):

\[sin\theta = -\frac{8}{17}\] Then sub it in the formula. Since \(cos2x\) is a double angle formulae

OpenStudy (mimi_x3):

Find \(cos2\theta\) right?